OpenStudy (dark_wolf):

Solve for x. 1+|2+x|=91+|2+x|=9 A. x = 6 or x=−10x=−10 B. x = 5 or ​ x=−9x=−9 ​ C. x = 7 or ​ x=−11x=−11 ​ D. x = 4 or ​ x=−8x=−8 ​

OpenStudy (dark_wolf):

@RedR0se03399 I have a question i cant solve :(

OpenStudy (redr0se03399):

wassup?

OpenStudy (dark_wolf):

not much

OpenStudy (dark_wolf):

you there?

563blackghost (563blackghost):

It looks like your question got repeated....

OpenStudy (redr0se03399):

sorry @Dark_Wolf im kinda dealing with some stuffs rn, ill answer your question as soon as i can

OpenStudy (sshayer):

\[1+\left| 2+x \right|=9,\left| 2+x \right|=9-1=8,2+x=\pm 8 \] find x